For the below Cash Flow, find the total PW value using 10% interest rate 


years cost $
0     3,255
1     1,000.00
2                 -  
3     4,112
4                 -  
5     4,000.00
6     1,000.00
Answer #1

$$ \begin{aligned} \frac{\varepsilon A}{(1+i)^{n}} \\ &=3225+\frac{1000}{(1+i)}+\frac{4112}{(1+1)^{3}}+\frac{4000}{(1+i)^{5}}+\frac{1500}{(1+i)^{6}} \\ &=3225+\frac{1000}{1.1}+\frac{4112}{1.1^{3}}+\frac{4060}{1.1^{5}}+\frac{1500}{1.1^{6}} \\ &=\$ 10271.66 \end{aligned} $$

answered by: mathboy